Grade 12Physical Chemistry

The average K.E of an ideal gas In calories per mole is approximately equal to:

Rea Shejith
9 Years agoGrade 12

2 Answers

Vikas TU
9 Years ago
The average K.E of an ideal gas In calories per mole is approximately equal to:
 
Every atoms is associated with an avergae kinetic energy as => 3KT/2
From the PV = nRT
n = PV/RT.
Prafulla patil
7 Years ago
We know that, 
K.E=3/2nRT
R=2 cal mol^-
n=1
Therefore, 
K.E=3/2*2*T
K.E=3T
Thai is, 3 times a absolute temp